OpenStudy (anonymous):

the length of a rectangle is 6 meters more than its width. If the area is 40 square meters, what are the length and width?

OpenStudy (anonymous):

\[x+\]\[w+6 = length\] \[So width is 4 meters and length is 10 meters\]

OpenStudy (anonymous):

let the width be x and length be x+6 area=l*w 40=(x+6)(x) x^2+6x=40 x^2+6x-40=0 now solve

OpenStudy (anonymous):

x(x+6)=40 ==> x^2 +6x-40=0 ===>(x-4)(x+10) = 10 ===> x=4 ( forget the negative solution) ===> width = 4 ===> length =10

OpenStudy (anonymous):

width=4m, length=10m